the average of four numbers is 638.4 the first three numbers are 216,159.8,301.25. what is the fourth number?
Answer by stanbon(57387)

the average of four numbers is 638.4 the first three numbers are 216,159.8,301.25. what is the fourth number?
----
Let the 4th number be "x":
(216+159.8+301.25 + x)/4 = 638.4
-----
677.05 + x = 4*638.4
x = 1876.55
===========
Cheers,
Stan H.